OpenStudy (anonymous):

Find the range of the given data set -4, -3, -1, -1, 0, 1

OpenStudy (anonymous):

You first need to put it in order. Least to greatest.

OpenStudy (anonymous):

0,-1,-1,-3,-4,1

OpenStudy (anonymous):

Now, the range means subtracting the big number and the lowest number. What is the biggest and smallest number that is given?

OpenStudy (anonymous):

whats the answer tho, if i knew how i would of done it so whats the answer

OpenStudy (anonymous):

I can't give the answer to you. I'll help you out. The highest number is 1 and the lowest number is 0. 1-0 = ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Then that's your answer.

OpenStudy (anonymous):

but it aint the answer

OpenStudy (anonymous):

Which of the following best describes a chart of the number of tomatoes on a tomato plant each morning? discrete and linear discrete and non-linear continuous and linear continuous and non-linear

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

What are your answer choices?

OpenStudy (anonymous):

Ik your answer but can you give me mine first

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

-5 -3 3 5

OpenStudy (anonymous):

can you help me?

OpenStudy (anonymous):

Which of the following best describes a chart of the number of tomatoes on a tomato plant each morning? discrete and linear discrete and non-linear continuous and linear continuous and non-linear

OpenStudy (anonymous):

nah back up they helping me

OpenStudy (anonymous):

@lucastann the answer would be A.

OpenStudy (anonymous):

Dude please i will fan you and give you a medal

OpenStudy (anonymous):

ok thanks vsfreakz

OpenStudy (anonymous):

you gave him the answer but not me K

OpenStudy (anonymous):

@alondra_ hmm. I'm not sure what I did wrong 'cause the rules of the range is subtract big number to the smallest number.

OpenStudy (anonymous):

how about this

OpenStudy (anonymous):

Make a table using points along the graphed linear function. Then, determine the correct function. Line plotted on the coordinate plane that intersects the x-axis at zero and intersects the y-axis at zero. The line passes point 2 comma 1. y is two times x y is two more than x y is one-half x y is one-half more than x

OpenStudy (anonymous):

Oh give me a minute. I think I figured out what I did wrong.

OpenStudy (johnweldon1993):

*The smallest number is NOT 0...*

OpenStudy (anonymous):



OpenStudy (anonymous):

wassup what the answer

OpenStudy (anonymous):

Would the lowest number be -1? @johnweldon1993 Just to be sure I'm right.

OpenStudy (anonymous):

@lucastann please post this on a new forum question.

OpenStudy (anonymous):

how i am new here

OpenStudy (anonymous):

Look on your left side "Ask a question..."

OpenStudy (anonymous):

i don't see that

OpenStudy (johnweldon1993):

No the lowest number would be the -4... In order from LEAST to GREATEST we have -4, -3, -1, -1, 0, 1

OpenStudy (anonymous):

1 - (-4) = 1 + 4 = 5

OpenStudy (anonymous):

@johnweldon1993 thank you! I needed that help.